OpenStudy (anonymous):

note: e = is an element of... For m e N(natural numbers), let Cm = {x e R(real numbers) | m - 1 <= x^2 < m}. Is C = {Cm | m e N(natural numbers)} a partition of R(real numbers)?

OpenStudy (anonymous):

Quote: `jflush` ``` note: e = is an element of... ``` How cute!

OpenStudy (anonymous):

Try using `\[ \]` and `\( \)` around math.

OpenStudy (anonymous):

For \( m \in \mathbb{N} \) let \(C_m = \{x \in \mathbb{R} | m - 1 \le x^2 \lt m\}\). Is \(C = \{C_m | m \in \mathbb{N}\}\) a partition of \(\mathbb{R}\)? \]

OpenStudy (anonymous):

Okay so the real question is: 1) Is does every real number belong to some set in our partition? \(\forall r\in \mathbb{R} \;\exists m\in \mathbb{N}: \;r\in C_m\)? 2) Does no real number belong to more than one set in our partition? \(\forall n,k\in \mathbb{N}\neg \exists r: r\in C_m\wedge r\in C_k\wedge m=k\)